# Deep Submodular Peripteral Networks

### v2

@@ Appendix B Submodular Functions and Weighted Matroid Rank
Submodular set functions effectively capture notions of diversity and representativeness. Continuing on from §1, it is known that they are useful in machine learning in the context of data subset selection [Mirzasoleiman et al., 2020, Killamsetty et al., 2021b, a, Wei et al., 2015], feature selection [Liu et al., 2013], summarization [Lin and Bilmes, 2011], active learning [Wei et al., 2015, Kothawade et al., 2021], experimental design [Bhatt et al., 2024], more recently in reinforcement learning [Prajapat et al., 2023] and several other areas that involve some form of discrete optimization. Submodular functions also appear naturally in many other areas, such as theoretical computer science and statistical physics [Tohidi et al., 2020, Bilmes, 2022].
Submodular functions are also of interest since they enjoy properties sufficient for efficient optimization. For a ground set of size $n$, they lie in a cone of dimension $2^{n}$ in $\mathbb{R}^{2^{n}}$, but can still be minimized exactly without constraints in polynomial time [Fujishige, 2005]. Even though submodular function maximization is NP-hard, it offers a fast constant factor approximation algorithm, for instance in the cardinality-constrained case, greedy results in $1-1/e$ guarantee for polymatroid functions [Nemhauser et al., 1978, Minoux, 2005]. Similarly one can give guarantees for other constrained cases such as knapsack constraint or more complicated independence or matroid constraints [Buchbinder et al., 2014, Calinescu et al., 2011, Lee et al., 2009, Iyer and Bilmes, 2013, Iyer et al., 2013].
@@
We use the (conditional) gain notation $f(v|X)\triangleq f(X\cup\{v\})-f(X)$ regularly in the paper. We also assume also that all submodular functions are monotone non-decreasing (or just monotone), which means that for any $X\subseteq Y$ we have that $f(X)\leq f(Y)$ and non-negative, meaning that $0\leq f(X)$ for any $X\subseteq V$. We also assume $f$ is normalized so that at the emptyset $f(\emptyset)=0$. Such normalized monotone non-decreasing submodular functions necessarily are always non-negative, and are commonly referred to as polymatroid functions [Bilmes, 2022, Cunningham, 1983, 1985, Lovász, 1980].
@@
For completeness, we note that an equivalent definition of submodularity would say that for any two sets $A,B\subseteq V$, we have that $f(A)+f(B)\geq f(A\cup B)+f(A\cap B)$. It is not hard to show that this and Inequality (8) are identical [Bilmes, 2022, Fujishige, 2005].
@@
Matroids [Lawler, 1976, Oxley, 2011, Schrijver, 2003] are useful algebraic structures that are strongly related to submodular functions [Fujishige, 2005]. Given a finite ground set $V$ of size $|V|=n$, a matroid $\mathcal{M}$ is a set system that consists of the pair $(V,\mathcal{I})$ where $V$ is said ground set and =$\mathcal{I}=\{I_{1},I_{2},\dots\}$ is a set of subsets of $V$, meaning $I_{i}\subseteq V$ for all $i$. To define matroid $\mathcal{M}$ we say $\mathcal{M}=(V,\mathcal{I})$. The subsets $\mathcal{I}$ are known as the independent sets (meaning, for any $I\in\mathcal{I}$, $I$ is said to be independent) and they have certain properties that make this the set system interesting and useful. Formally, a matroid is defined as follows:
@@
A matroid rank functions can be significantly extended when it is associated with a non-negative vector $m\in\mathbb{R}^{|V|}_{+}$ of length $|V|=n$. With such a vector, one can define a weighted matroid rank function:
@@
    Consider a uniform matroid $M=(V,\mathcal{I})$ where $\mathcal{I}=2^{V}$. This means that all $2^{n}$ subsets of $V$ are independent in the matroid. Then given non-negative vector $m\in\mathbb{R}_{+}^{|V|}$, the weighted matroid rank function has the form $\text{rank}_{\mathcal{M},m}(A)=\sum_{a\in A}m(A)$, thus the weighted matroid rank function is just the modular function $m$.
@@
Another interesting (and apparently previously unreported) property of weighted matroid rank functions is that they are always permutation invariant in the same sense as a deep set function [Zaheer et al., 2017] is permutation invariant. Firstly, we restate the definition of permutation invariance from [Zaheer et al., 2017].
Definition 6 (Permutation Invariance [Zaheer et al., 2017]).
@@
Appendix C Primer on Deep Submodular Functions
Stobbe and Krause [2010] proposed a scalable algorithm for minimizing a class of submodular functions known as decomposable functions. In general, given a collection of modular functions, $\{m_{i}\}^{N}_{i=1}$ where each $m_{i}:V\to\mathbb{R}_{+}$ and $\phi$, a monotone non-decreasing, normalized $\phi(0)=0$ and non-negative concave function,
@@ Appendix C Primer on Deep Submodular Functions
is said to be a “decomposible” function.
In machine learning, objects of interest are often represented by featurizing them into a finite-dimensional vector space. If $U=\{u_{1},u_{2},\ldots,u_{|U|}\}$ indexes the dimensions, then for any example $v\in V$ the modular function $m_{U}(v)=\left(m_{u_{1}}(v),m_{u_{2}}(v),\ldots,m_{u_{|U|}}(v)\right)^{T}$ represents the features of $v$. $m_{u}(v)$ can be interpreted in several ways such as the degree of presence of $u-th$ concept in $v$ similar to concept bottleneck models [Koh et al., 2020] or even the count of an n-gram feature [Wei et al., 2014, Lin and Bilmes, 2011, 2012]. With this nomenclature, we re-define decomposable submodular functions as feature-based functions that has the form:
@@
One issue with the feature-based function happens when there is redundancy between two features $u_{i}$ and $u_{j}$. For instance, if the features are a count of n-grams, and since higher-order n-grams encompass lower-order n-grams, having too many of a certain high-order n-gram would also mean having many of its lower-order n-grams, leading to redundancy (or correlation) in what the features represent. As eq. 13 suggests, feature-based functions treat each feature independently. This motivates a flexible class of submodular functions, known as Deep submodular functions or DSFs [Bilmes and Bai, 2017] which mitigates the problem of redundancy by an additional concave composition followed by weighted aggregation which allows for interaction between sets of redundant features as follows:
@@
Here, $w_{s,u}$ and $\omega_{s}$ are non-negative mixture weights. This naturally leads to a representation resembling neural networks when we iteratively apply a concave function followed by linear aggregation, giving rise to Deep Submodular Functions. $\forall s\in S$,  Bilmes and Bai [2017] define $\sum_{u\in U}w_{s,u}\phi_{u}\left(m_{u}(A)\right)$ as meta-features, and in general one can stack more layers of concave composition followed by affine transformations, leading to meta-meta features, and so on. Let $L$ be the total number of layers in a DSF; each meta-feature at layer $\ell$ of this network is a submodular function, we index them using $U^{(\ell)}$, which means in the example in equation (14) $U=U^{(0)}$ and $S=U^{(1)}$.
@@
here, $\cdot$" denotes matrix multiplication. To maintain submodularity, all weight matrices must be non-negative. A DSF can incorporate skip connections, and in general, a directed acyclic graph topology can be used to represent the DSF architecture with the help of recursive definitions (refer to section 4.1 of Bilmes and Bai [2017]).
@@
Lastly, Bilmes and Bai [2017] showed that DSFs strictly generalize feature-based functions, and adding additional layers strictly increases the representation capacity, therefore suggesting the utility of additional layers!
@@ Appendix D Weighted Matroid Rank Aggregation for DSFs
In theorem 12 we show that $f_{\mathcal{M}}$ is a permutation invariant submodular set function. To show that $f_{\mathcal{M}}$ is submodular we first state some properties that will be useful. We re-visit the definitions from  Bilmes and Bai [2017] for superdifferential and antitone superdifferential of a concave function which we will use to provide sufficient conditions that will help to show $f$ is submodular.
@@
The proof can be done following the similar steps as in Lemma 5.13, Corollary 5.13.1 and Lemma 5.14 from Bilmes and Bai [2017]. ∎
We now re-state Theorem 5.12 in Bilmes and Bai [2017] as the follows.
@@
The proof is the same as for Theorem 5.12 in Bilmes and Bai [2017]. ∎
@@
The output of each pillar is a $d$-dimensional non-negative vector. The aggregation via any weighted matroid rank of the pillar outputs for any set $A\subseteq V$ is also a $d$-dimensional non-negative vector and moreover, each element of this resulting aggregated vector is itself a polymatroid function thanks to the fact that a weighted matroid rank function is a polymatroid function. When we compose a vector of polymatroid functions with a DSF, this preserves polymatroidality by Theorem 5.12 in Bilmes and Bai [2017], and thus submodularity. The theorem is hence proven. ∎
@@ Appendix E Further discussion on (E,M) pair sampling and dataset construction strategies
Moreover, a set $E$ might be seen as diverse in some cases and homogeneous in other cases. For instance, a set $E$ containing elements from two distinct classes is heterogeneous compared to a set $M$ containing all elements from one class. However, the same set $E$ is homogeneous when compared against a set $E^{\prime}$ containing all distinct classes, so in this latter case the $(E,M)$ pair would be $(E^{\prime},E)$. The oracle query offers a GPC of the pair $(E,M)$ providing a score $\text{\emph{Score}}({A},{B})\in\mathbb{R}$ depending on which of $E$ and $M$ is more diverse. The pairwise comparison aspect of GPC is one of its powers since we are not asking for absolute scores from the oracle but for comparisons between pairs. This is one of the benefits of pairwise comparison models when querying humans as well (as discussed in the field of psychometrics [Nunnally and Bernstein, 1994] and this benefit extends to the GPC case as well [Lingel et al., 2022].
@@
Appendix F Analysis of Peripteral Loss
In addition to the motivation for our loss function provided in section 3 here we provide a discussion on its gradient and curvature and how the hyperparameters govern the curvature of the provided loss. First, we re-state the mother loss, plot its value for different values of margin, and then follow by equation for its gradient and double-derivative. In the below $\sigma(x)$ refers to sigmoid function, that is, $\sigma(x)=\frac{1}{1-e^{-x}}$ and $\operatorname{sgn}(x)$ refers to the sign of $x$.

### v3

@@ Appendix B Submodular Functions and Weighted Matroid Rank
Submodular set functions effectively capture notions of diversity and representativeness. Continuing on from §1, it is known that they are useful in machine learning in the context of data subset selection [69, 45, 44, 106], feature selection [62], summarization [59], active learning [106, 49], experimental design [11], more recently in reinforcement learning [79] and several other areas that involve some form of discrete optimization. Submodular functions also appear naturally in many other areas, such as theoretical computer science and statistical physics [101, 12].
Submodular functions are also of interest since they enjoy properties sufficient for efficient optimization. For a ground set of size $n$, they lie in a cone of dimension $2^{n}$ in $\mathbb{R}^{2^{n}}$, but can still be minimized exactly without constraints in polynomial time [28]. Even though submodular function maximization is NP-hard, it offers a fast constant factor approximation algorithm, for instance in the cardinality-constrained case, greedy results in $1-1/e$ guarantee for polymatroid functions [73, 68]. Similarly one can give guarantees for other constrained cases such as knapsack constraint or more complicated independence or matroid constraints [15, 16, 54, 40, 39].
@@
We use the (conditional) gain notation $f(v|X)\triangleq f(X\cup\{v\})-f(X)$ regularly in the paper. We also assume also that all submodular functions are monotone non-decreasing (or just monotone), which means that for any $X\subseteq Y$ we have that $f(X)\leq f(Y)$ and non-negative, meaning that $0\leq f(X)$ for any $X\subseteq V$. We also assume $f$ is normalized so that at the emptyset $f(\emptyset)=0$. Such normalized monotone non-decreasing submodular functions necessarily are always non-negative, and are commonly referred to as polymatroid functions [12, 20, 21, 64].
@@
For completeness, we note that an equivalent definition of submodularity would say that for any two sets $A,B\subseteq V$, we have that $f(A)+f(B)\geq f(A\cup B)+f(A\cap B)$. It is not hard to show that this and Inequality (6) are identical [12, 28].
@@
Matroids [53, 77, 88] are useful algebraic structures that are strongly related to submodular functions [28]. Given a finite ground set $V$ of size $|V|=n$, a matroid $\mathcal{M}$ is a set system that consists of the pair $(V,\mathcal{I})$ where $V$ is said ground set and =$\mathcal{I}=\{I_{1},I_{2},\dots\}$ is a set of subsets of $V$, meaning $I_{i}\subseteq V$ for all $i$. To define matroid $\mathcal{M}$ we say $\mathcal{M}=(V,\mathcal{I})$. The subsets $\mathcal{I}$ are known as the independent sets (meaning, for any $I\in\mathcal{I}$, $I$ is said to be independent) and they have certain properties that make this the set system interesting and useful. Formally, a matroid is defined as follows:
@@
A matroid rank functions can be significantly extended when it is associated with a non-negative vector $m\in\mathbb{R}^{|V|}_{+}$ of length $|V|=n$. With such a vector, one can define a weighted matroid rank function that returns the maximum weight of any base of a set $A$:
@@
    Consider a free matroid $M=(V,\mathcal{I})$ where $\mathcal{I}=2^{V}$. This means that all $2^{n}$ subsets of $V$ are independent in the matroid. Then given non-negative vector $m\in\mathbb{R}_{+}^{|V|}$, the weighted matroid rank function has the form $\text{rank}_{\mathcal{M},m}(A)=\sum_{a\in A}m(a)$, thus the weighted matroid rank function is just the modular function $m$.
@@
Another interesting (and apparently previously unreported) property of weighted matroid rank functions is that they are always permutation invariant in the same sense as a deep set function [110] is permutation invariant. Firstly, we restate the definition of permutation invariance from [110].
Definition 7 (Permutation Invariance [110]).
@@
Appendix C Primer on and Novelty over Deep Submodular Functions
In this section, we offer a brief background on deep submodular functions (DSFs) and demonstrate how our results in the present paper are not just incremental over this previous work — rather the present paper offers significant novelty over, as far as we can tell from an extensive literature search, previously published papers.
[96] proposed a scalable algorithm for minimizing a class of submodular functions known as decomposable functions. In general, given a collection of modular functions, $\{m_{i}\}^{N}_{i=1}$ where each $m_{i}:V\to\mathbb{R}_{+}$ and $\phi$, a monotone non-decreasing, normalized $\phi(0)=0$ and non-negative concave function,
@@ Appendix C Primer on and Novelty over Deep Submodular Functions
is said to be a “decomposable” function.
In machine learning, objects of interest are often represented by featurizing them into a finite-dimensional vector space. If $U=\{u_{1},u_{2},\ldots,u_{|U|}\}$ indexes the dimensions, then for any example $v\in V$ the modular function $m_{U}(v)=\left(m_{u_{1}}(v),m_{u_{2}}(v),\ldots,m_{u_{|U|}}(v)\right)^{T}$ represents the features of $v$. $m_{u}(v)$ can be interpreted in several ways such as the degree of presence of $u-th$ concept in $v$ similar to concept bottleneck models [48] or even the count of an n-gram feature [107, 59, 60]. With this nomenclature, we re-define decomposable submodular functions as feature-based functions that has the form:
@@
One issue with the feature-based function happens when there is redundancy between two features $u_{i}$ and $u_{j}$. For instance, if the features are a count of n-grams, and since higher-order n-grams encompass lower-order n-grams, having too many of a certain high-order n-gram would also mean having many of its lower-order n-grams, leading to redundancy (or correlation) in what the features represent. As eq. 11 suggests, feature-based functions treat each feature independently. This motivates a flexible class of submodular functions, known as Deep submodular functions or DSFs [13] which mitigates the problem of redundancy by an additional concave composition followed by weighted aggregation which allows for interaction between sets of redundant features as follows:
@@
Here, $w_{s,u}$ and $\omega_{s}$ are non-negative mixture weights. This naturally leads to a representation resembling neural networks when we iteratively apply a concave function followed by linear aggregation, giving rise to Deep Submodular Functions. $\forall s\in S$,  [13] define $\sum_{u\in U}w_{s,u}\phi_{u}\left(m_{u}(A)\right)$ as meta-features, and in general one can stack more layers of concave composition followed by affine transformations, leading to meta-meta features, and so on. Let $L$ be the total number of layers in a DSF; each meta-feature at layer $\ell$ of this network is a submodular function, we index them using $U^{(\ell)}$, which means in the example in equation (12) $U=U^{(0)}$ and $S=U^{(1)}$.
@@
here, $\cdot$" denotes matrix multiplication. To maintain submodularity, all weight matrices must be non-negative. A DSF can incorporate skip connections, and in general, a directed acyclic graph topology can be used to represent the DSF architecture with the help of recursive definitions (refer to section 4.1 of [13]).
@@
Lastly, [13] showed that DSFs strictly generalize feature-based functions, and adding additional layers strictly increases the representation capacity, therefore suggesting the utility of additional layers.
It is important to realize that our present work on DSPNs and GPC-based learning is not just an incremental improvement to DSFs, but rather offers a new class of submodular function and an entirely new way to learn the submodular function. There are two pre-existing papers on DSFs, this paper [23] from 2016 as well as a more theoretical paper [13] from 2017. The first paper [23] introduces DSFs and uses a max-margin based learning algorithm that was introduced by [60] in their “Submodular Shells” paper but applies that learning algorithm to DSFs. Max-margin learning of submodular functions requires a dataset of highly valued subsets according to some unknown submodular function to be learnt (e.g., each subset can be thought of as a good summary) and the goal of max-margin learning in this case is to learn a submodular function so that it scores those subsets highly relative to other subsets. Max-margin learning does not care at all about the rest of the valuations of the submodular function (e.g., a max-margin learnt submodular function is under no obligation to perform well for problems other than submodular maximization, and is under no obligations to preserve the rank valuations of a given teacher submodular function).
Our present paper is entirely different from these two papers. While our DSPN “roof” structure uses the deep concave function from DSFs, there is no commonality beyond this. In the present (DSPN) paper, we have introduced a new parametric submodular function class (DSPNs) that overcomes the only known weakness of DSFs (namely, DSFs inability to express certain graphic matroid rank functions, as proven in [13]). DSPNs do not have this weakness thanks to the submodularity-preserving permutation-invariant aggregation stage (see Lemma 1 and Corollaries 3 and 4) which immediate means that DSPNs can represent any graphic matroid rank function. The expressivity improvement of DSPNs over DSFs means that it is now an open theoretical problem if DSPNs can represent all polymatroid functions or not. Considering the narrow range of submodular functions that DSFs cannot represent, we believe that it is likely that DSPNs indeed can represent any polymatroid function, but we do not have a proof of this at the present time.
We have also, in the present paper, introduced a new learning paradigm (the peripteral loss) for learning DSPNs that does not suffer from the issues that the max-margin learning approach suffers from. In essence, our learning is a form of distillation from an expensive (possibly unoptimizable) target oracle down to a parametric and computationally efficient student (a DSPN). That is, with a peripteral loss, the graded comparison signal $\Delta$ can be any positive/negative value, which could come from a difference in submodular valuations, or could come from a non-submodular function, or could come from a human evaluator (we explain this in detail in Section 1). I.e., the learnt submodular function can, via minimizing the peripteral loss, scale well for all values of any target function (e.g., low valued sets should be given a low value by the learnt function, and high valued set should be given a high value by the learnt function, in order for the peripteral loss to be small). In the present study, we choose $\Delta$ as the difference between expensive-to-optimize facility location function evaluations since it is a widely and successfully used submodular function but that suffers from computational scalability issues when used for optimization (e.g., submodular max) over large datasets.
Another critical novel component of our paper over [23, 13] is that, to the best of our knowledge, our peripteral loss approach is the first to establish a connection between: (I) contrastive learning approaches in representation learning (see Section A.1); (II) pairwise comparison scaling methods widely used in RLHF (and LLM learning, as discussed in the paper, lines 55–81); and (III) submodular learning. We establish connections between these three approaches while at the same time extending regular pairwise comparison approaches to the graded pairwise comparison approach (GPC), which is further novelty. While we do not study contrastive learning or RLHF in our paper, we believe that a broad impact of our work is that our peripteral loss for GPC will be useful for both higher-order contrastive learning methods and for RLHF, but we leave that to future studies as we focus on the submodular learning problem in the present paper (and, as we discuss in our introduction, we believe there is too little attention in the community paid to practical strategies for identifying a good submodular functions for a problem).
Lastly, the experimental results in [23] consider only very small-scale (and entirely offline) summarization experiments (summarizing 100 images down to 10) or synthetic data experiments. By contrast, we consider experiments on a much larger scale (130K images) and apply it towards both offline and online summarization problems in the context of choosing a good set of training data subset points that have not been previously explored — in other words, we instantiate the resulting DSPNs on entirely held-out data, so not only do DSPNs generalize to different sets and set sizes of a given ground set $V$ but DSPNs generalize to entirely different ground sets $V^{\prime}$ (meaning $V^{\prime}\cap V=\emptyset$).
So in sum, the present paper offers enormous novelty and difference over the two previous DSF papers [23, 13].
@@ Appendix D Weighted Matroid Rank Aggregation for DSFs
In theorem 13 we show that $f_{\mathcal{M}}$ is a permutation invariant submodular set function. To show that $f_{\mathcal{M}}$ is submodular we first state some properties that will be useful. We re-visit the definitions from  [13] for superdifferential and antitone superdifferential of a concave function which we will use to provide sufficient conditions that will help to show $f$ is submodular.
@@
The proof can be done following the similar steps as in Lemma 5.13, Corollary 5.13.1 and Lemma 5.14 from [13]. ∎
We now re-state Theorem 5.12 in [13] as the follows.
@@
The proof is the same as for Theorem 5.12 in [13]. ∎
@@
The output of each pillar is a $d$-dimensional non-negative vector. The aggregation via any weighted matroid rank of the pillar outputs for any set $A\subseteq V$ is also a $d$-dimensional non-negative vector and moreover, each element of this resulting aggregated vector is itself a polymatroid function thanks to the fact that a weighted matroid rank function is a polymatroid function. When we compose a vector of polymatroid functions with a DSF, this preserves polymatroidality by Theorem 5.12 in [13], and thus submodularity. The theorem is hence proven. ∎
@@ Appendix E Further discussion on (E,M) pair sampling and dataset construction strategies
Moreover, a set $E$ might be seen as diverse in some cases and homogeneous in other cases. For instance, a set $E$ containing elements from two distinct classes is heterogeneous compared to a set $M$ containing all elements from one class. However, the same set $E$ is homogeneous when compared against a set $E^{\prime}$ containing all distinct classes, so in this latter case the $(E,M)$ pair would be $(E^{\prime},E)$. The oracle query offers a GPC of the pair $(E,M)$ providing a score $\text{\emph{Score}}({A},{B})\in\mathbb{R}$ depending on which of $E$ and $M$ is more diverse. The pairwise comparison aspect of GPC is one of its powers since we are not asking for absolute scores from the oracle but for comparisons between pairs. This is one of the benefits of pairwise comparison models when querying humans as well (as discussed in the field of psychometrics [74] and this benefit extends to the GPC case as well [61].
@@
Appendix F Analysis and Further Description of Peripteral Loss
In addition to the motivation for our loss function provided in section 3 here we provide a discussion on its gradient and curvature and how the hyperparameters govern the curvature of the provided loss. First, we re-state the mother loss, plot its value for different values of margin, and then follow by equation for its gradient and double-derivative. In the below $\sigma(x)$ refers to sigmoid function, that is, $\sigma(x)=\frac{1}{1+e^{-x}}$ and $\operatorname{sgn}(x)$ refers to the sign of $x$.